# HybridNorm: Towards Stable and Efficient Transformer Training via Hybrid Normalization

### v1

@@ 4 Experiments
Table 1: Downstream evaluation results of 1.2B dense models with Pre-Norm, HybridNorm, and HybridNorm under 1T training tokens.
Methods BasicArithmetic HellaSwag SciQ ARC-C ARC-E PIQA OpenbookQA COPA Avg.$\uparrow$
@@
In this section, we demonstrate the effectiveness of HybridNorm through extensive experiments in LLMs.
4.1 Experiment Settings
Baseline. We evaluate HybridNorm across two series of models: dense models and Mixture of Experts (MoE) models. The dense models include two scales: 550M and 1B, with the latter containing approximately 1.27 billion parameters and utilizing an architecture similar to Llama 3.2 (Dubey et al., 2024). All analytical experiments are conducted on the 550M dense models. For the MoE model, we use the OLMoE framework (Muennighoff et al., 2025), which activates 1.3B parameters out of a total of 6.9B parameters (MoE-1B-7B). Both models are trained from scratch on the OLMoE Mix dataset  [^1] OLMoE Mix is available at https://huggingface.co/datasets/allenai/OLMoE-mix-0924. (Muennighoff et al., 2025).
Model Configuration. The 550M dense model has a model dimension of 1536, an FFN dimension of 4096, and utilizes 16 attention heads with 4 key/value heads per attention head. The 1.2B model features a larger model dimension of 2048 and an FFN dimension of 9192, with 32 attention heads and 8 key/value heads per attention head. The MoE-1B-7B model employs 16 attention heads, a model dimension of 2048, and an FFN dimension of 1024. Notably, it features 8 experts out of 64, providing a more fine-grained distribution of computational resources. All models consist of 16 layers and are trained with a consistent context length of 4096. More details can be found in Appendix B.
Hyperparameters. Model weights are initialized using Megatron initialization (Shoeybi et al., 2019) (See Section 4.4 for more details). For the optimization, we apply the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. All models are trained on sequences of 4096 tokens. For the dense model, we set the initial learning rate to 3e-4, decaying to 3e-5 using a cosine scheduler. The MoE model starts with a learning rate of 4e-4, decaying according to a cosine schedule. We summarize the hyperparameters in Table 7.
Evaluation Metrics. To evaluate the performance of LLMs with HybridNorm, we employ a diverse set of open benchmarks, including ARC-Easy (ARC-E) (Clark et al., 2018), ARC-Challenge (ARC-C) (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), CoQA (Reddy et al., 2019), Winogrande (Sakaguchi et al., 2021), MMLU (Hendrycks et al., 2021), BoolQ (Clark et al., 2019), COPA (Gordon et al., 2012), CSQA (Talmor et al., 2019), OBQA (Mihaylov et al., 2018), and SocialIQA (Sap et al., 2019). We leverage the LM Eval Harness (Gao et al., 2023) for standardized performance evaluation.
4.2 Main Results on Dense Models
We evaluate the performance of HybridNorm and HybridNorm on 1.2B dense transformer models. Figure 1 compares the training dynamics of dense models with different normalization methods. As shown in the figure, models with HybridNorm and HybridNorm exhibit consistently lower training loss and validation perplexity throughout training compared to Pre-Norm, highlighting their effectiveness in enhancing training stability and convergence.
Table 1 presents the downstream evaluation results across a diverse set of benchmarks. HybridNorm consistently outperforms Pre-Norm in most tasks, achieving the highest average score. Notably, it demonstrates substantial improvements in tasks such as BasicArithmetic (+3.11), HellaSwag (+1.71), and COPA (+3.78), indicating enhanced generalization and robustness. These results underscore the scalability of HybridNorm in larger transformer models, further validating its effectiveness in improving both training stability and downstream performance. More results can be found in Figure 7.
@@ 4.2 Main Results on Dense Models
Table 2: Downstream evaluation results of MoE-1B-7B with Pre-Norm and HybridNorm under 500B training tokens.
Methods HellaSwag ARC-C ARC-E PIQA WinoGrande OpenbookQA BoolQ COPA Avg.$\uparrow$
@@
4.3 Main Results on MoE Models
For MoE models, we conduct experiments on MoE-1B-7B with 8 experts selected from a pool of 64. Figure 3 presents the training dynamics of MoE models under different normalization strategies. Throughout the training, HybridNorm consistently achieves lower training loss and validation perplexity compared to Pre-Norm. These findings indicate that HybridNorm effectively alleviates optimization difficulties in large-scale MoE models, resulting in more stable training and enhanced downstream performance.
Further, as shown in Table 2, HybridNorm consistently outperforms Pre-Norm across various downstream tasks, achieving the highest average score. Notably, it demonstrates significant improvements in ARC-C (+2.35), ARC-E (+2.40), and OpenbookQA (+0.81), highlighting its ability to enhance generalization across diverse benchmarks.
@@ 4.4 Ablation Studies
Pre-Norm Normal 2.75 20.29
Depth-Scaled 2.76 20.49
Megatron 2.76 20.44
HybridNorm Normal 2.76 20.44
Depth-Scaled 2.76 20.40
Megatron 2.74 20.00
Initialization. To evaluate the sensitivity of Pre-Norm and HybridNorm to initialization schemes, we conduct ablation studies comparing three widely used initialization strategies: Normal initialization (Nguyen & Salazar, 2019), Depth-Scaled initialization (Zhang et al., 2019; Gururangan et al., 2023), and Megatron initialization (Shoeybi et al., 2019). Normal initialization initializes all weights of linear layers using a truncated normal distribution with mean zero and standard deviation $1/\sqrt{2.5d}$, where $d$ is the hidden dimension. Depth-Scaled initialization and Megatron initialization introduce scaling factors to stabilize training in deep architectures. Specifically, Depth-Scaled initialization scales down the output projections of the attention and FFN by a factor of $\sqrt{2l}$, where $l$ is the layer index. In contrast, Megatron initialization scales down these projections by $\sqrt{2L}$, where $L$ is the total number of layers, mitigating gradient variance accumulation in very deep transformers.
As shown in Table 3, Pre-Norm and HybridNorm exhibit sensitivity across different initialization methods, achieving the lowest training loss and perplexity under Normal initialization and Megatron initialization, respectively. Therefore, we set the default initialization method for Pre-Norm to Normal initialization and for HybridNorm to Megatron initialization in all experiments, respectively, which ensures that even under settings that may be more favorable to baseline models, the superiority of our approach is effectively demonstrated.
@@
Normalization Position. We investigate the impact of the position of normalization layers within the transformer block. First, we examine the effect of varying the placement of QKV normalization (e.g., normalization setting in attention). We extend the normalization setting by considering not only the Query (Q), Key (K), and Value (V) components but also the Context (C), which refers to the output of the attention mechanism. For instance, QKVC-Norm applies normalization to all four components: Query, Key, Value, and Context, while KV-Norm and KC-Norm focus on the normalization of the Key-Value and Key-Context pairs, respectively. QKVC-Post refers to transformer blocks that employ QKVC-Norm in the MHA while using Post-Norm in the FFN. Second, we explore the effect of integrating QKV-Norm into different transformer architectures. For instance, Pre-QKV-Pre refers to a configuration where QKV-Norm is applied with Pre-Norm in the MHA layer, while the FFN layer utilizes Post-Norm. Other configurations follow similar definitions. Finally, we compare various hybrid combinations of Pre-Norm and Post-Norm. Pre-Post refers to transformer blocks that apply Pre-Norm in the MHA and Post-Norm in the FFN, whereas Post-Pre adopts the opposite configuration. Mathematical formulas for methods mentioned above can be found in Appendix E.
As shown in Table 4, HybridNorm (also called QKV-Post) and its variant HybridNorm consistently surpass other configurations. Notably, HybridNorm achieves the lowest training loss and perplexity while attaining the highest accuracy on HellaSwag. Specifically, by comparing HybridNorm with the first block in Table 4, we find that QKV-Norm is the most effective normalization setting in the attention. Similarly, comparing HybridNorm with the second block, we observe that combining QKV-Norm with Post-Norm in the FFN yields superior performance. From the third block, one can see that the Pre-Post configuration indeed leads to improved performance, while replacing Pre-Norm in the MHA with QKV-Norm to form HybridNorm further enhances performance, achieving the best results.
@@
Special Treatment of First Block. For the special treatment of the first block, we test different architectures, such as adding a normalization layer after embedding (call EmbedNorm) and armed the first block with QKV-norm and Pre-Norm in FNN (call First-QKV-Pre), which formulations are:
$Y^{0}$ $=\mathrm{MHA}_{QKV}(X^{0})+X^{0},$
$X^{1}$ $=\mathrm{FFN}(\mathrm{Norm}(Y^{0}))+Y^{0}.$
@@
Figure 5: Gradient norm comparison of Pre-Norm, Post-Norm, and HybridNorm at different training steps.
@@
4.5 Gradient Analysis
To gain deeper insights into the stability improvements introduced by HybridNorm, we analyze gradient norms across training iterations. As shown in Figure 5, we compare the gradient norms of Pre-Norm, Post-Norm, and HybridNorm at steps 1 and 100. The results indicate that Pre-Norm tends to exhibit gradient explosion in deeper models, while Post-Norm suffers from vanishing gradients, both of which hinder effective optimization. In contrast, HybridNorm maintains a well-balanced gradient flow throughout training, effectively mitigating these issues. An intuitive understanding is that Pre-Norm tends to amplify gradients, while Post-Norm diminishes them. HybridNorm alternates between these two normalization strategies, leading to more stable gradient propagation during backpropagation and effectively preventing gradient explosion or vanishing. This balanced gradient propagation contributes to smoother optimization dynamics and faster convergence, further reinforcing the effectiveness of HybridNorm in stabilizing transformer training.
Figure 6: Scaling law curves of Pre-Norm and HybridNorm.
@@ 4.7 Deeper Models
To further evaluate the robustness of HybridNorm and HybridNorm in deeper architectures, we conduct experiments on transformers with depths ranging from 16 to 29 layers while maintaining a comparable parameter budget. This setup allows for a fair comparison of different normalization strategies in deep transformer architectures. As shown in Table 5, both HybridNorm and HybridNorm consistently outperform Pre-Norm and Post-Norm across various depths, demonstrating their effectiveness in stabilizing deep model training.
A particularly striking observation is that Post-Norm fails to converge at 29 layers, reinforcing its well-documented instability in deeper architectures. In contrast, HybridNorm and HybridNorm not only ensure stable training across all depths but also achieve significantly lower training loss and perplexity on both the C4 and Pile datasets. These improvements indicate that HybridNorm-based normalization strategies mitigate optimization difficulties that commonly arise in deep transformers. Furthermore, HybridNorm achieves the highest accuracy on challenging downstream benchmarks such as HellaSwag and PIQA, suggesting that its benefits extend beyond mere training stability to enhanced generalization on real-world tasks. These results provide strong empirical evidence that HybridNorm-based normalization schemes enable deeper transformer training while preserving superior optimization efficiency and downstream task performance.

### v3

@@ 5 Experiments
5.1 Experiment Settings
Baseline
We evaluate HybridNorm across two series of models: dense models and Mixture of Experts (MoE) models. The dense models include two scales: 550M and 1B, with the latter containing approximately 1.27 billion parameters and utilizing an architecture similar to LLaMA 3.2 [11]. All analytical experiments are conducted on the 550M dense models. For the MoE model, we use the OLMoE framework [24], which activates 1.3B parameters out of a total of 6.9B parameters (MoE-1B-7B). Both models are trained from scratch on the OLMoE Mix dataset [24].
Model Configuration
The 550M dense model has a model dimension of 1536, an FFN dimension of 4096, and utilizes 16 attention heads with 4 key/value heads per attention head. The 1.2B model features a larger model dimension of 2048 and an FFN dimension of 9192, with 32 attention heads and 8 key/value heads per attention head. The MoE-1B-7B model employs 16 attention heads, a model dimension of 2048, and an FFN dimension of 1024. Notably, it features 8 experts out of 64, providing a more fine-grained distribution of computational resources. All models consist of 16 layers and are trained with a consistent context length of 4096. More details can be found in Appendix 9.
Hyperparameters
Model weights are initialized using Megatron initialization [33] (See Section 5.3 for more details). For the optimization, we apply the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. All models are trained on sequences of 4096 tokens. For the dense model, we set the learning rate to 3e-4, decaying to 3e-5 using a cosine scheduler. The MoE model starts with a learning rate of 4e-4, decaying with a cosine schedule. We summarize the hyperparameters in Table 7.
Evaluation Metrics
To evaluate the performance of LLMs with HybridNorm, we employ a diverse set of open benchmarks, including ARC-Easy (ARC-E) [5], ARC-Challenge (ARC-C) [5], HellaSwag [46], PIQA [2], SciQ [41], CoQA [29], Winogrande [31], MMLU [15], BoolQ [4], COPA [13], CSQA [36], OBQA [23], and SocialIQA [32]. We leverage the LM Eval Harness [12] for standardized performance evaluation.
Figure 3: Training dynamics for 1.2B dense models with Pre-Norm, HybridNorm and HybridNorm under 1T training tokens. We present the training loss, validation loss, and downstream performance on HellaSwag and ARC-Easy, demonstrating that HybridNorm achieves superior performance.
Table 1: Downstream evaluation results of 1.2B dense models with Pre-Norm, HybridNorm, and HybridNorm under 1T training tokens. OQA refers to OpenbookQA.
Methods BasicArithmetic HellaSwag SciQ ARC-C ARC-E PIQA OQA COPA Avg.$\uparrow$
@@ 5.1 Experiment Settings
5.2 Main Results
Dense Models
We evaluate the performance of HybridNorm and HybridNorm on 1.2B dense transformer models. Figure 3 compares the training dynamics of dense models with different normalization methods. As shown in the figure, models with HybridNorm and HybridNorm exhibit consistently lower training loss and validation perplexity throughout training compared to Pre-Norm, highlighting their effectiveness in enhancing training stability and convergence. Table 1 presents the downstream evaluation results. HybridNorm consistently outperforms Pre-Norm in most tasks, achieving the highest average score. Notably, it demonstrates substantial improvements in tasks such as BasicArithmetic (+3.11), HellaSwag (+1.71), and COPA (+3.78), indicating enhanced generalization and robustness. These results underscore the scalability of HybridNorm in larger transformer models, further validating its effectiveness in improving both training stability and downstream performance. More results can be found in Figure 9. In Appendix 10.1, we present additional comparisons with other approaches, such as Post-Norm and Mix-LN. We further provide analyses of signal propagation and entropy dynamics across different methods, as detailed in Appendix 10.2 & 10.3.
MoE Models
For MoE models, we conduct experiments on MoE-1B-7B with 8 experts selected from a pool of 64. Figure 4 presents the training dynamics of MoE models under different normalization strategies. Throughout the training, HybridNorm consistently achieves lower training loss and validation perplexity compared to Pre-Norm. These findings indicate that HybridNorm effectively alleviates optimization difficulties in large-scale MoE models, resulting in more stable training and enhanced downstream performance. Further, as shown in Table 2, HybridNorm consistently outperforms Pre-Norm across various downstream tasks, achieving the highest average score. Notably, it demonstrates significant improvements in ARC-C (+2.35), ARC-E (+2.40), and OpenbookQA (+0.81), highlighting its ability to enhance generalization across diverse benchmarks.
@@ 5.2 Main Results
Table 2: Downstream evaluation results of MoE-1B-7B with Pre-Norm and HybridNorm under 500B training tokens. OQA refers to OpenbookQA.
Methods HellaSwag ARC-C ARC-E PIQA WinoGrande OQA BoolQ COPA Avg.$\uparrow$
@@ 5.3 Ablation Studies
Pre-Norm Normal 2.75 20.29 HybridNorm Normal 2.76 20.44
Depth-Scaled 2.76 20.49 Depth-Scaled 2.76 20.40
Megatron 2.76 20.44 Megatron 2.74 20.00
Initialization
To evaluate the sensitivity of Pre-Norm and HybridNorm to initialization schemes, we conduct ablation studies comparing three widely used initialization strategies: Normal initialization [25], Depth-Scaled initialization [48, 14], and Megatron initialization [33]. Normal initialization initializes all weights of linear layers using a truncated normal distribution with mean zero and standard deviation $1/\sqrt{2.5d}$, where $d$ is the hidden dimension. Depth-Scaled initialization and Megatron initialization introduce scaling factors to stabilize training in deep architectures. Specifically, Depth-Scaled initialization scales down the output projections of the attention and FFN by a factor of $\sqrt{2l}$, where $l$ is the layer index. In contrast, Megatron initialization scales down these projections by $\sqrt{2L}$, where $L$ is the total number of layers, mitigating gradient variance accumulation in very deep transformers. As shown in Table 3, Pre-Norm and HybridNorm exhibit sensitivity across different initialization methods, achieving the lowest training loss and perplexity under Normal initialization and Megatron initialization, respectively. Therefore, we set the default initialization method for Pre-Norm to Normal initialization and for HybridNorm to Megatron initialization in all experiments, respectively, which ensures that even under settings that may be more favorable to baseline models, the superiority of our approach is effectively demonstrated.
@@
Normalization Position
We investigate the impact of the position of normalization layers within the transformer block. First, we examine the effect of varying the placement of QKV normalization (e.g., normalization setting in attention). We extend the normalization setting by considering not only the Query (Q), Key (K), and Value (V) components but also the Context (C), which refers to the output of the attention mechanism. For instance, QKVC-Norm applies normalization to all four components: Query, Key, Value, and Context, while KV-Norm and KC-Norm focus on the normalization of the Key-Value and Key-Context pairs, respectively. QKVC-Post refers to transformer blocks that employ QKVC-Norm in the MHA while using Post-Norm in the FFN. Second, we explore the effect of integrating QKV-Norm into different transformer architectures. For instance, Pre-QKV-Post refers to a configuration where QKV-Norm is applied with Pre-Norm in the MHA layer, while the FFN layer utilizes Post-Norm. Other configurations follow similar definitions. Finally, we compare various hybrid combinations of Pre-Norm and Post-Norm. Pre-Post refers to transformer blocks that apply Pre-Norm in the MHA and Post-Norm in the FFN, whereas Post-Pre adopts the opposite configuration. Mathematical formulas for methods mentioned above can be found in Appendix 12.
As shown in Table 4, HybridNorm (a.k.a. QKV-Post) and its variant HybridNorm consistently surpass other methods. Notably, HybridNorm achieves the lowest training loss and perplexity while attaining the highest accuracy on HellaSwag. Specifically, by comparing HybridNorm with the left first block in Table 4, we find that QKV-Norm is the most effective normalization. Similarly, comparing HybridNorm with the left second block, we observe that combining QKV-Norm with Post-Norm in the FFN yields superior performance. From the right table, one can see that the Pre-Post configuration indeed leads to improved performance, while replacing Pre-Norm in the MHA with QKV-Norm to form HybridNorm further enhances performance, achieving the best results.
@@
Figure 6: Scaling law curves of Pre-Norm and HybridNorm.
Special Treatment of First Block
For the special treatment of the first block, we test different architectures, such as adding a normalization layer after embedding (call EmbedNorm) and armed the first block with QKV-norm and Pre-Norm in FFN (call First-QKV-Pre), which formulations are:
$Y^{0}$ $=\mathrm{MHA}_{QKV}(X^{0})+X^{0},\quad X^{1}=\mathrm{FFN}(\mathrm{Norm}(Y^{0}))+Y^{0}.$ (10)
@@ 5.5 Deeper Models
To further evaluate the robustness of HybridNorm and HybridNorm in deeper architectures, we conduct experiments on transformers with depths ranging from 16 to 29 layers while maintaining a comparable parameter budget. This setup allows for a fair comparison of different normalization strategies in deep transformer architectures. As shown in Table 5, both HybridNorm and HybridNorm consistently outperform Pre-Norm and Post-Norm across various depths, demonstrating their effectiveness in stabilizing deep model training. A particularly striking observation is that Post-Norm fails to converge at 29 layers, reinforcing its well-documented instability in deeper architectures. In contrast, HybridNorm and HybridNorm not only ensure stable training across all depths but also achieve significantly lower training loss and perplexity on both the C4 and Pile datasets. These improvements indicate that HybridNorm-based normalization strategies mitigate optimization difficulties that commonly arise in deep transformers. Furthermore, HybridNorm achieves the highest accuracy on challenging downstream benchmarks such as HellaSwag and PIQA, suggesting that its benefits extend beyond mere training stability to enhanced generalization on real-world tasks. These results provide strong empirical evidence that HybridNorm-based normalization schemes enable deeper transformer training while preserving superior optimization efficiency and downstream task performance.